# Adversarial Robust Decision Transformer: Enhancing Robustness of RvS via Minimax Returns-to-go

### v1

@@ 4.1 Full Data Coverage Setting
The efficacy of our solution is first evaluated on three short-horizon sequential games with adaptive adversary: (1) a single-stage game, (2) an adversarial, single-stage bandit environment depicting a Gambling round (Paster et al., 2022), and (3) a multi-stage game. These are depicted in Figure 1, Figure 6, and Figure 8, respectively. The collected data consists of $10^{5}$ trajectories, encompassing all possible trajectories. The online policies for data collection employed by the protagonist and adversary were both uniformly random.
@@ 4.2 Discrete Game with Partial Data Coverage
DT 0.18 (0.09) -0.28 (0.05) -0.57 (0.06) -1.00 (0.00)
ESPER 0.44 (0.09) 0.00 (0.10) -0.42 (0.05) -0.98 (0.01)
ARDT (ours) 0.55 (0.11) 0.11 (0.22) 0.02 (0.09) -1.00 (0.00)
@@
Connect Four is a discrete sequential two-player game with deterministic dynamics (Paster et al., 2022), which can also be viewed as a decision-making problem when facing an adaptive adversary. There is no intermediate reward; rather, there is a terminal reward of either $-1$, 00, or $1$, meaning lose, tie or win for the first-moving player, respectively. Therefore we fix the target return in Connect Four as $1$. We fix the protagonist to be the first-move player. This game has a maximum horizon of $22$ steps. Since we have a solver of this game, we manage to collect the data with $\epsilon$-greedy policy. We define the optimality of an $\epsilon$-greedy policy of protagonist or adversary as $(1-\epsilon)\cdot 100\%$. Each dataset includes a total of $10^{6}$ number of steps for variable-length trajectories.
@@ 4.3 Continuous Adversarial Environments
Datasets DT ESPER ARDT (ours)
Hopper-lr 112.10 (63.40) 66.80 (5.50) 477.90 (219.90)
Hopper-mr 330.20 (286.80) 103.20 (11.70) 482.20 (83.70)
Hopper-hr 298.62 (319.72) 63.00 (11.50) 331.69 (187.18)
Walker2D-lr 429.60 (31.90) 380.60 (156.90) 405.60 (39.80)
Walker2D-mr 391.00 (20.60) 426.90 (29.70) 508.40 (61.80)
Walker2D-hr 413.22 (54.69) 502.60 (31.30) 492.63 (109.28)
Halfcheetah-lr 1416.00 (48.50) 1691.80 (63.00) 1509.20 (33.80)
Halfcheetah-mr 1229.30 (171.10) 1725.60 (111.80) 1594.60 (137.50)
Halfcheetah-hr 1493.03 (32.87) 1699.80 (32.80) 1765.71 (150.63)
@@
In this section, we test the algorithms on a specific instance of Adversarial Reinforcement Learning problems, namely Noisy Action Robust MDP (NR-MDP), which was introduced by  Tessler et al. (2019). This includes three continuous control problems of OpenAI Gym MuJoCo, Hopper, Walker2D and Halfcheetah  (Todorov et al., 2012) involving the adversarial action noise. In NR-MDP, the protagonist and adversary act jointly. Specifically, they exact a total force over one or more parts of the body of the agent in a Mujoco environment, corresponding to the weighted sum of their individual forces (adversary’s weight $\delta$). Despite no longer being in the discrete setting, the return minimax expectile regression in ARDT is implemented with the raw continuous states and actions input, without any discretization.
To collect the data, we first trained Action Robust Reinforcement Learning (ARRL) online policies on NR-MDP tasks  (Tessler et al., 2019) to create the online protagonist policy and adversary policy for evaluation. Subsequently, the pre-collected data consist of the replayed trajectories removing the ones with low cumulative rewards. We collect the suboptimal trajectories with the $\epsilon$-greedy of the saved ARRL agents, i.e. online protagonist and adversary policies. We then mix them with the pre-collected datasets to create the final training datasets. The proportion of random trajectories to pre-collected trajectories are $0.01:1$, $0.1:1$ and $1:1$, respectively, for datasets categorised as being low, medium, and high randomness. At test time, we evaluate each of our trained ARDT protagonist policies against a population of $8$ adversarial online policies, each across $5$ seeds. We sweep over the same set of target returns for all algorithms, and pick the best results.
@@ 5 Conclusions and Limitations
This paper introduces a worst-case-aware training algorithm designed to improve the adversarial robustness of the Decision Transformer. By relabeling trajectories with the estimated in-sample minimax returns-to-go through expectile regression, our algorithm is demonstrated to be robust against adversaries more powerful than the behavior ones, which existing DT methods cannot achieve. In experiments, our method consistently exhibits superior worst-case performance against adversarial attacks in both gaming environments and continuous control settings.
@@ 6 Acknowledgments
Ilija Bogunovic was supported by the EPSRC New Investigator Award EP/X03917X/1; the Engineering and Physical Sciences Research Council EP/S021566/1; and Google Research Scholar award. Xiaohang Tang was supported by the Engineering and Physical Sciences Research Council [grant number EP/T517793/1, EP/W524335/1]. The authors would like to thank the Department of Statistical Science, in particular Chakkapas Visavakul, for co-ordinating the computer resources, and the Department of Electronic and Electrical Engineering at University College London for providing the computer clusters.
Supplementary Material
@@ Appendix B Environments
Multi-Stage Game.
Apart from using a single-stage game as a motivating example, we also conducted experiments on a multi-stage game to highlight the importance of multi-step minimax instead of a single-step minimum by an adversary. This is because the optimal adversary will also consider the optimal protagonist strategy in the subsequent states. A multi-step minimax return is more accurate in describing the value of states and sub-trajectories.
Figure 6: Multi-stage Game. LHS is the tree representation of the problem, where P and A are decision-maker (protagonist) and adversary, respectively. RHS is the same game tree where we attach the learned minimax return and the action probabilities of the well-trained ARDT of a single run when conditioning on the target return $7$ to the adversarial node and the branches, respectively. For example, the highlighted node has minimax return-to-go $4.9$. $0.99$ on the branch leads to the highlited node, which indicates the action probability of taking $a_{0}$ at the initial state. Thick lines represent the optimal adversarial actions.
Gambling.
Gambling  [Paster et al., 2022] is originally a simple stochastic bandit problem with three arms and horizon one. When taking the first arm, the reward will be either $5$ or $-15$ with $50\%$ probability each. Similarly, when taking the second arm, the reward will be $1$ or $-6$ with equal probabilities. The third arm has deterministic reward $1$. In this paper, we extend it to an adversarial setting by considering an adversary controlling the outcomes, shown in Figure 8.
@@
Connect Four.
Connect Four has a long horizon of $22$ in the worst case. Figure 8 shows an example of this game’s dynamics. The outcomes only consist of win, tie, and lose, leading to the reward for the protagonist $1$, 00, and $-1$. We leverage the solver in https://github.com/PascalPons/connect4 to generate a stochastic policy for data collection.
@@ Appendix C Implementation Details
Our implementation is based on the implementation of Decision Transformer [Chen et al., 2021], and ESPER [Paster et al., 2022]. Our data collection and part of the evaluation against online adversaries are based on the implementation in Noisy Action Robust Reinforcement Learning [Tessler et al., 2019].
The hyperparameters of the Minimax Expectile Regression are in Table 3. In the practical version of ARDT’s minimax regression, we adopt a leaf regularization to ensure the accuracy of predicting outcomes at the terminal state (leaf node). We record our training iterations of minimax regression in the hyperparameters table. The number of training iterations in full coverage examples and Connect Four are larger than twice the maximum trajectory length, guaranteeing the minimax return propagated from the leaves to every node. The predicted values are relatively accurate according to the prediction in Figure 6. While in the MuJoCo environment, the trajectory length is too long for minimax regression to make full propagation, partial propagate already brings robustness based on the MuJoCo results in Section 4.3.
Decision Transformer is trained with cross-entropy loss in the discrete environment and MSE in the continuous environment, as in Decision Transformer [Chen et al., 2021]. During action inference at test time, all DT-based methods are prompted with a desired initial target return and state to generate the strategy $\pi$. Subsequently, the trajectory is generated in an auto-regressive manner, utilizing both the actions sampled by the protagonist’s strategy $\pi$ and the adversarial strategy $\bar{\pi}$, along with the transitions and rewards provided by the online environments, till the environment terminates or the maximum trajectory length is reached for evaluation. Notably, the target return will be subtracted from the observed immediate reward once received from the environment.
Notably, given that the rewards and states tokens encompass sufficient information about the adversary’s action, we train the Transformer model with sequences without adversarial tokens, as it ensures that the model capacity remains consistent across algorithms. As usual, vanilla DT’s are conditioned on the observed returns-to-go, contaminated by the adversarial actions and states. Similarly, ESPER policy is conditional on the expected returns-to-go and states. Since the adversary only influences the transition and rewards, ESPER treats the adversary implicitly as part of the environment that causes stochasticity and addresses it with expected return as the information function.
@@
Table 5: Data profile of Connect Four. The tuples represent the optimal percentages of protagonist and adversary.
Appendix D Comparison to Other Methods
Tabular value-based works have been studied to solve zero-sum games offline by leveraging pessimism for out-of-sample estimation [Cui and Du, 2022, Zhong et al., 2022]. In contrast, we did not incorporate pessimism due to the difficulty of tuning its associated hyperparameter. We employ expectile regression to approximate the in-sample minimax return, inspired by Implicit Q Learning [Kostrikov et al., 2021] to approximate the Bellman optimal $Q$-values leveraging expectile regression (ER). ER is efficient since it doesn’t rely on querying the actions during minimax return estimation, thus reducing complexity. Learning value functions this way, and learning strategy from the data is much less complex. Thus, learning values and extracting policy from them is reasonable. In addition, training Transformer based on relabeled returns-to-go is similar to doing policy extraction from the learned values (returns-to-go) [Neumann and Peters, 2008]. Transformers demonstrating strong reasoning in sequence modeling should be exploited to improve reinforcement learning.
@@ Appendix E Data and Computing Resources
We publish our datasets of Connect Four and Mujoco data collected by executing Noisy Action Robust Reinforcement Learning and save the data on a Google drive: https://drive.google.com/drive/folders/1fd9RNmntL0p86ywEcRkbIKZbKTgJ6BY1?usp=sharing. There are no data access restrictions. The Mujoco data profiles are in Table 5 and 4. The MuJoCo data has a $1000$ number of trajectories, each with $1000$ steps of interactions. The Connect Four datasets also have a $10^{6}$ number of interaction steps, where each trajectory has a length of at most $22$.
We conducted experiments on GPUs: a GeForce RTX 2080 Ti with memory 11GB and an NVIDIA A100 with memory 80 GB. The memory for the entire computing is around 100GB when parallelizing the running over random seeds in the Connect Four game and MuJoCo.

### v2

@@ 4.1 Full Data Coverage Setting
The efficacy of our solution is first evaluated on three short-horizon sequential games with adaptive adversary: (1) a single-stage game, (2) an adversarial, single-stage bandit environment depicting a Gambling round [32], and (3) a multi-stage game. These are depicted in Figure 1, Figure 7, and Figure 9, respectively. The collected data consists of $10^{5}$ trajectories, encompassing all possible trajectories. The online policies for data collection employed by the protagonist and adversary were both uniformly random.
@@ 4.2 Discrete Game with Partial Data Coverage
DT 0.18 (0.09) -0.28 (0.05) -0.57 (0.06) -1 (0.0)
ESPER 0.44 (0.09) 0 (0.1) -0.42 (0.05) -0.98 (0.01)
ARDT (ours) 0.55 (0.11) 0.11 (0.22) 0.02 (0.09) -1 (0.0)
@@
Connect Four is a discrete sequential two-player game with deterministic dynamics [32], which can also be viewed as a decision-making problem when facing an adaptive adversary. There is no intermediate reward; rather, there is a terminal reward of either $-1$, 00, or $1$, meaning lose, tie or win for the first-moving player, respectively. Therefore we fix the target return in Connect Four as $1$. We fix the protagonist to be the first-move player. This game has a maximum horizon of $22$ steps. Since we have a solver of this game, we manage to collect the data with $\epsilon$-greedy policy. We define the optimality of an $\epsilon$-greedy policy of protagonist or adversary as $(1-\epsilon)\cdot 100\%$. Each dataset includes a total of $10^{6}$ number of steps for variable-length trajectories.
@@ 4.3 Continuous Adversarial Environments
Datasets DT ESPER ARDT
Hopper-lr 112.1 (63.4) 66.8 (5.5) 477.9 (219.9)
Hopper-mr 330.2 (286.8) 103.2 (11.7) 482.2 (83.7)
Hopper-hr 298.62 (319.72) 63.0 (11.5) 331.69 (187.18)
Walker2D-lr 429.6 (31.9) 380.6 (156.9) 405.6 (39.8)
Walker2D-mr 391.0 (20.6) 426.9 (29.7) 508.4 (61.8)
Walker2D-hr 413.22 (54.69) 502.6 (31.3) 492.63 (109.28)
Halfcheetah-lr 1416.0 (48.5) 1691.8 (63.0) 1509.2 (33.8)
Halfcheetah-mr 1229.3 (171.1) 1725.6 (111.8) 1594.6 (137.5)
Halfcheetah-hr 1493.03 (32.87) 1699.8 (32.8) 1765.71 (150.63)
@@
In this section, we test the algorithms on a specific instance of Adversarial Reinforcement Learning problems, namely Noisy Action Robust MDP (NR-MDP), which was introduced by  Tessler et al. [42]. This includes three continuous control problems of OpenAI Gym MuJoCo, Hopper, Walker2D and Halfcheetah  [43] involving the adversarial action noise. In NR-MDP, the protagonist and adversary act jointly. Specifically, they exact a total force over one or more parts of the body of the agent in a Mujoco environment, corresponding to the weighted sum of their individual forces (adversary’s weight $\delta$). Despite no longer being in the discrete setting, the return minimax expectile regression in ARDT is implemented with the raw continuous states and actions input, without any discretization.
To collect the data, we first trained Action Robust Reinforcement Learning (ARRL) online policies on NR-MDP tasks  [42] to create the online protagonist policy and adversary policy for evaluation. Subsequently, the pre-collected data consist of the replayed trajectories removing the ones with low cumulative rewards. We collect the suboptimal trajectories with the $\epsilon$-greedy of the saved ARRL agents, i.e. online protagonist and adversary policies. We then mix them with the pre-collected datasets to create the final training datasets. The proportion of random trajectories to pre-collected trajectories are $0.01:1$, $0.1:1$ and $1:1$, respectively, for datasets categorised as being low, medium, and high randomness. At test time, we evaluate each of our trained ARDT protagonist policies against a population of $8$ adversarial online policies, each across $5$ seeds. We sweep over the same set of target returns for all algorithms, and pick the best results.
@@
4.4 Ablation Study
Finally, we offer abalation study on the expectile level $\alpha$ in discrete environments. Specifically, we varied $\alpha$ for relabeling in Single-stage Game, Multi-stage Game and Connect Four. Then we tested against the optimal adversary in the first two environments and an $\epsilon$-greedy adversary ($\epsilon=0.5$) for the third environment.
According to the first two environments, we can confirm a pattern consistent with our theory that smaller alpha leads to more robustness. According to the Equation (6)-(9) in the paper, it should be that when the $\alpha$ is closer to 0, our estimate of minimax return is more accurate, and thus our algorithm is more robust. When the alpha is increased to 0.5, our method is converted to Expected-Return-Conditioned Decision Transformer (e.g. ESPER) since the expectile loss in this case reduces to Mean Square Error. This can be confirmed by comparing the results of Single-stage game and Multi-stage game in Figure 6 and 3.
In the third environment, Connect Four, the performance initially increases as decreases, but eventually drops. This is due to the weak stochastic adversary. When the value is too small, the algorithm becomes overly conservative. However, ARDT still outperforms DT significantly. Additionally, it also implies that we can tune the parameter alpha to adjust the level of conservativeness
Figure 6: Ablation study on expectile level $\alpha$ over $10$ random seeds. In connect four, we set the target return to be $1$, and the adversary to be $\epsilon$-greedy ($\epsilon=0.5$), i.e. taking uniform random policy with probability $0.5$. On Connect Four, ARDT with $\alpha=0.01$ acts too conservative.
@@ 5 Conclusions and Limitations
This paper introduces a worst-case-aware training algorithm designed to improve the adversarial robustness of the Decision Transformer, namely Adversarially Robust Decision Transformer (ARDT). By relabeling trajectories with the estimated in-sample minimax returns-to-go through expectile regression, our algorithm is demonstrated to be robust against adversaries more powerful than the behavior ones, which existing DT methods cannot achieve. In experiments, our method consistently exhibits superior worst-case performance against adversarial attacks in both gaming environments and continuous control settings.
@@ 6 Acknowledgments
Ilija Bogunovic was supported by the EPSRC New Investigator Award EP/X03917X/1; the Engineering and Physical Sciences Research Council EP/S021566/1; and Google Research Scholar award. Xiaohang Tang was supported by the Engineering and Physical Sciences Research Council [grant number EP/T517793/1, EP/W524335/1]. The authors would like to thank the Department of Statistical Science, in particular Chakkapas Visavakul, for co-ordinating the computer resources, and the Department of Electronic and Electrical Engineering, and Department of Computer Science at University College London for providing the computer clusters.
Appendix
@@ Appendix B Environments
Multi-Stage Game
Apart from single-stage game as motivating example, we also conduct experiments also on a multi-stage game in order to highlight the importance of multi-step minimax instead of a single-step minimum by adversary. This is because the optimal adversary will also consider the optimal protagonist strategy in the subsequent states. Considering a multi-step minimax return is more accurate to describe the value of states and sub-trajectories.
Figure 7: Multi-stage Game. LHS is the tree representation of the problem, where P and A are decision maker (protagonist) and adversary, respectively. RHS is the same game tree where we attach the learned minimax return and the action probabilities of the well-trained ARDT of a single run when conditioning on the target return $7$ to the adversarial node and the branches, respectively. For example, the highlighted node has minimax return-to-go $4.9$. $0.99$ on the branch lead to the highlited node indicates the action probability of taking $a_{0}$ at the initial state. Thick lines represent the optimal adversarial actions.
Gambling
Gambling  [32] is originally a simple stochastic bandit problem with three arms and horizon one. When taking the first arm, the reward will be either $5$ or $-15$ with $50\%$ probability each. Similarly, when taking the second arm, the reward will be $1$ or $-6$ with equal probabilities. The third arm has deterministic reward $1$. In this paper, we extend it to adversarial setting by considering an adversary controlling the outcomes, shown in Figure 9.
@@
Connect Four
Connect Four has a long horizon $22$ in the worst case. An example of the dynamics of this game is displayed in Figure 9. The outcomes only consist of win, tie, and lose, leading to the reward for the protagonist $1$, 00, and $-1$. We leverage the solver in https://github.com/PascalPons/connect4 to generate stochastic policy for data collection.
@@ Appendix C Implementation Details
Our implementation is based on the implementation of Decision Transformer [2], and ESPER [32]. Our data collection and part of the evaluation against online adversaries are based on the implementation in Noisy Action Robust Reinforcement Learning [42].
The hyperparameters of the Minimax Expectile Regression are in Table 3. In practical version of ARDT’s minimax regression, we adopt a leaf regularization to ensure the accuracy of predicting outcomes at terminal state (leaf node). We record our training iterations of minimax regression in the hyperparameters table. The number of training iterations in full coverage examples and Connect Four are larger than twice of the maximum trajectory length, which guarantees the minimax return propagated from the leaves to every node. According to the prediction in Figure 7, the predicted values are relatively accurate. While in MuJoCo environment, the trajectory length are too long for minimax regression to make full propagation, partial propagate already brings robustness based on the MuJoCo results in Section 4.3.
Decision Transformer is trained with cross entroy loss in discrete environment and MSE in continuous environment, as in Decision Transformer [2]. During action inference at test time, all DT-based methods are prompted with a desired initial target return and state to generate the strategy $\pi$. Subsequently, the trajectory is generated in an auto-regressive manner, utilizing both the actions sampled by protagonist’s strategy $\pi$ and the adversarial strategy $\bar{\pi}$, along with the transitions and rewards provided by the online environments, till the environment terminates or the maximum trajectory length is reached for evaluation. Notably, the target return will be subtracted by the observed immediate reward once received any from the environment.
Notably, given that the rewards and states tokens encompass sufficient information about the adversary’s action, we train the Transformer model with sequences without adversarial tokens, as it ensures that the model capacity remains consistent across algorithms. As usual, vanilla DT’s is conditioned on the observed returns-to-go, which are contaminated by the adversarial actions, and states. Similarly, ESPER policy is conditional on the expected returns-to-go and states. Since the adversary only influences the transition and rewards, ESPER treats the adversary implicitly part of the environment who causes stochasticity, and addresses it with expected return as the information function.
@@
Table 5: Data profile of Connect Four. The tuples represent the optimalities percentages of protagonist and adversary.
Appendix D Additional Related Works
D.1 Algorithmically Related Works
Tabular value-based works have been studied to solving zero-sum games offline by leveraging pessimism for out-of-sample estimation [4, 56]. In contrast, we did not incorporate pessimism due to the difficulty of tuning its associated hyperparameter. We employ expectile regression to approximate the in-sample minimax return, inspired by the way of Implicit Q Learning [16] to approximate the Bellman optimal $Q$-values leveraging expectile regression (ER). ER is efficient since it doesn’t rely on querying the actions during minimax return estimation, thus reducing complexity. Learning value functions in this way has much less complexity to learn strategy from the data directly. Thus, the protocol of learning values and then extracting policy from them is reasonable. In addition, training Transformer based on relabeled returns-to-go is similar to do policy extraction from the learned values (returns-to-go) [28]. Transformer which demonstrates the strong reasoning in sequence modeling, should be exploited for improving Reinforcement Learning.
D.2 Related Adversarial Attacks
Within the broad setting of Robust Reinforcement Learning [27], our experiments focus on action-robustness.We conducted experiments on Noisy Action Robust-MDP as a representative environment, following the previous works [7, 42, 14, 34]. According to the different types of attacks proposed in [25, 35], our adversarial perturbation can be viewed as Man In The Middle (MITM) attack formulation proposed in [35], specifically actions attack, as noted in the last paragraph of the Related Work section of [35]. The actions attack we considered in our work can influence all elements in RL from the learner’s perspective, including the next state observation, immediate reward, and dynamics, thus potentially causing all types of perturbation introduced in [25, 35]. Therefore, it is reasonable to select actions attack as a representative problem, although our architecture can be extended well to other formulations of robustness to the attacks.
D.3 Related Robust RL Methods
Robust Reinforcement Learning can be roughly categorized into addressing training-time robustness to poisoning attacks, and test-time robustness to evasion attacks [49]. Poisoning attacks are defined as the manipulation of elements in the training trajectory, including state, actions, and rewards in the data or online environment [25, 35]. The adversary in this paper is similar to the Man In The Middle (MITM) attack formulation in [35]. Robust IQL [55] and RL agent with survival instinct [22] have demonstrated robustness to data poisoning in an offline setting. However, they only aim at the training-time robustness, while testing the algorithms without any attack. Our method is closer to ROLAH [7] and other related works [42, 14, 34], considering adversary appearing in both training-time and test-time serving evasion attacks. Despite the similarity,these methods are all online. Our method learns to achieve the robustness offline, which is more challenging due to potential distribution shift, and the lack of good behavior policy. In this paper, We didn’t select offline robust RL as baseline since we only focus on the Reinforcement Learning via Supervised Learning (RvS) methods. We study the robustness of Decision Transformer (DT), and thus provide more analysis on DT-based algorithms including vanila DT and ESPER in various offline learning settings and environments.
@@ Appendix E Data and Computing Resources
We publish our datasets along with the codebase via https://github.com/xiaohangt/ardt. There is no data access restrictions. The Mujoco data profiles are in Table 5 and 4. The MuJoCo data has $1000$ number of trajectories, each with $1000$ steps of interactions. The Connect Four datasets have also $10^{6}$ number of steps of interaction in total, where each trajectory has a length at most $22$.
We conduct experiments on GPUs: a GeForce RTX 2080 Ti with memory 11GB, and a NVIDIA A100 with memory 80GB. The memory for the entire computing is around 100GB when parallelizing the running over random seeds in Connect Four game, and MuJoCo.
@@ Appendix F Additional Experimental Results
ARDT can empirically address stochastic environments. Intuitively, this is because by tuning the expectile level , we can balance between expected return and minimax return, as discussed above, to achieve robustness against the distributional environmental changes.